# GETReason: Enhancing Image Context Extraction through Hierarchical Multi-Agent Reasoning

### v1

@@ Abstract
Shikhhar Siinghthanks:   Equal contribution.  Abhinav Rawat11footnotemark: 1  Vivek Gupta  Chitta Baral Arizona State University {ssiingh, arawat11, vgupt140, chitta}@asu.edu
@@ 1 Introduction
1.1 Motivation
The challenges above underscore why extracting event-specific context from public event images remains difficult. Existing approaches either limit themselves to shallow visual cues or depend on external retrieval, which risks hallucinations, misinformation, and fragmented context. True event understanding requires reasoning about temporal, spatial, and sociopolitical factors—often implicit rather than visually explicit.
@@ 1.1 Motivation
1.2 Main Contributions
Our approach is driven by the need for a systematic, reasoning-based method to extract three key components from an image: (1) geospatial location (e.g., "Indonesia"), (2) temporal information (e.g., "October 20, 2019"), and (3) event-specific details (e.g., "inauguration of President Joko Widodo’s second term amid public tensions"). To overcome prior limitations, we introduce a structured, multi-agentic framework where specialized agents collaboratively extract and validate contextual elements. Inspired by Ng et al. (2024), this architecture reduces errors and hallucinations through cross-agent validation, improving both precision and reliability. A thorough working example for the framework is explained in section Appendix under Working Example.
@@ 2 Dataset Description
The TARA dataset comprises of approximately 16,000 images associated with news articles, along with their corresponding time and location information, automatically extracted from the New York Times (NYT) [^1] https://developer.nytimes.com/docs/archive-product/1/overview. For our evaluation, we utilized the train set which comprises of 12,306 samples from this dataset, spanning from January 2010 to May 2021. Each image is accompanied by metadata detailing the time and location pertinent to the associated news article.
@@ 2.1 TARA*
To better suit our task, we augment TARA dataset with event and fine-grained temporal and geospatial information generated from the metadata present in the ground truth using a Vision Language Model (VLM), specifically, Gemini 1.5 Pro [^2] https://deepmind.google/.
@@ 5 Experiment
For GPT-4o-mini [^3] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, we utilized the batch API. While OpenAI’s models do not directly reveal public figures [^4] https://openai.com/policies/usage-policies/, they can ground them if their identities are provided externally in the prompt. To leverage this, we used the CelebrityFaces API from AWS Rekognition [^5] https://aws.amazon.com/rekognition/ to identify public figures in the dataset images with a confidence score of at least 99%. These identified public figures were then explicitly included in the prompts during the experiments to enable more precise entity recognition.
@@
5.2 Baselines
Reasoning-Based Approaches.
Object Detection-Based Approaches.
RAG-Based Approaches.
@@ 5.8 Effectiveness Analysis
Fundamental methodological differences
  • Hierarchical Multi-Agent Framework: Our method employs a hierarchical multi-agent framework, reducing cognitive load on the VLM, minimizing hallucinations, and enabling context resetting—unlike baselines that rely on single, monolithic prompts.
  • Structured Prompting: We use structured, role-specific prompts and controlled information flow, providing greater consistency and interpretability than baselines that depend on ad-hoc prompt engineering.
  • Schema-Driven Reasoning: By combining structured schema extraction with flexible reasoning, our approach achieves both precision and adaptability, addressing the inefficiencies and hallucinations often seen in less structured methods.
  • Systematic Cue Integration: Our framework explicitly handles temporal, geospatial, and event-specific cues, which are often overlooked or inconsistently extracted in prior methods (see Section 5.2).
  • Robust, Large-Scale Evaluation: GETReason is evaluated on over 17,000 examples, enabling a more robust and generalizable assessment than prior works limited to datasets of around 2,000 samples.
  • Diverse Baseline Limitations: As detailed in Section 5.2, baseline methods span object detection, shallow reasoning, and retrieval-based approaches—each introducing specific limitations in contextual depth, prompt precision, or overreliance on external sources.
@@ 6 Conclusion
The main limitation of this approach is the language barrier, as it relies on English-language datasets. Additionally, the TARA dataset’s extraction method results in many images being unrelated to their corresponding articles. Our analysis is based on a subset of TARA, whereas a more comprehensive evaluation could be conducted on the full 61k dataset or even larger alternatives like SHERLOCK Hessel, Hwang, Park, Zellers, Bhagavatula, Rohrbach, Saenko, and Choi (2022) and BreakingNews dataset Ramisa, Yan, Moreno-Noguer, and Mikolajczyk (2017) along with the comparison with newer studies like PuzzleGPT Ayyubi, Feng, Liu, Lin, Wang, and Chang (2025) to better assess data loss and the study’s limitations. Furthermore, as context extraction depends on model reasoning and Vision-Language Models (VLMs) are rapidly advancing, our findings may not fully reflect the latest developments. Although we used three competitive models, more robust reasoning based alternatives like OpenAI-o1666https://openai.com/o1/ , OpenAI-o3-mini [^7] https://openai.com/index/openai-o3-mini/ and DeepSeek-R1 Team (2025a) have since emerged.
@@ Acknowledgements
We thank the anonymous reviewers for their helpful feedback.
We also thank Zenia Shaikh (zshaikh8@asu.edu) for her valuable assistance and contributions to this work.
Our work is sponsored by the Army Research Office and is accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-19-1-2620.
Appendix
Individual Agentic Prompts:
Scene Graph Agent prompt.
You are an expert in analyzing and understanding visual images to generate detailed and structured scene graphs in JSON-formatted outputs. You perform the following two tasks sequentially: Scene Graph Generation: For every image provided, identify all entities, objects, attributes, and relationships in the scene. The graph must include all common entities, whether singular or plural, such as "car," "person," "crowd," "tree," "building," etc., as well as their properties and the spatial or contextual relationships among them. Your goal is to create a complete and accurate representation of all visible elements in the scene. Graph Enhancement: Refine the scene graph by identifying and replacing the generic descriptions of entities that correspond to public figures provided in the input list. For instance, if a "person" in the scene is identified as a specific public figure, replace "person" with their provided name. During this process, ensure that: All entities, whether identified or unidentified, remain in the graph. All relationships between objects and entities are preserved. Unidentified or unrecognized entities are left unchanged but still included in the graph. Your responses must be concise but thorough, breaking down the image into its components and interactions while ensuring that the scene graph remains fully intact, comprehensive, and user-friendly. Do not prune out entities or relationships during any part of the process. The objective is to provide users with a refined scene graph that includes all visible entities and incorporates specific public figure names where applicable, enhancing both clarity and usability.You may or may not be provided with the public figures present in the image. If present, ground them to their common nouns in their entities.You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Abstract Agent prompt.
@@
Derive an abstract_idea that captures the overall essence or message of the scene.
@@
You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA. Scene Graph: <generated scene graph>
Prompt Agent Prompt.
@@
Geospatial specialist - Analyzes locations from images and scene graphs with detailed reasoning. You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA. Scene Graph:<abstract added scene graph>
Geospatial Agent Direct Extraction prompt.
You are a custom agent designed to analyze an image and any accompanying contextual information to deduce geospatial details about the scene depicted. The details to determine are: - Country - State_or _Province - City Using the visual elements of the image and the provided contextual information, you must determine the best possible value for each of these fields. If you cannot deduce a particular field with reasonable confidence, output “NA” for that field. Rules: - You must provide exactly one value for each of the following fields: “Country,” “State_or _Province,” and “City.” You will return the output in the required response format. It is absolutely imperative that you return the JSON output.You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.", Scene Graph: <abstract added scene graph> Prompt: <prompt agent generated geospatial agent prompt>
Temporal Agent Direct Extraction prompt.
You are an agent designed to deduce the time of day, time period, or year depicted in an image based on its visual elements, the scene graph, and a descriptive prompt. . When analyzing an image, consider the following:
@@
- Best Guess: If conclusive reasoning is not possible, provide the most plausible estimate and explain your reasoning.You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA. Scene Graph: <abstract added scene graph> Prompt: <prompt agent generated temporal agent prompt>
Event Agent Direct Extraction prompt.
@@
3. Best Guess: Make logical assumptions if the inputs are inconclusive, ensuring relevance to the global context. Tailor responses to be clear, neutral, and informative, avoiding bias or unfounded speculation.You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA. Scene Graph: <abstract added scene graph> Prompt: <prompt agent generated event agent prompt>
Geospatial Agent Cross Extraction prompt.
You are a custom agent designed to analyze an image and any accompanying contextual information to deduce geospatial details about the scene depicted, paying special attention to the added event and temporal information along with the event’s background. It is important to note that any null, NA, unknown values in the provided are ignored. The details to determine are:
-͡ Country
-͡ State_or_Province
-͡ City
Using the visual elements of the image and the provided contextual information, you must determine the best possible value for each of these fields. If you cannot deduce a particular field with reasonable confidence, output “NA” for that field.
@@
-͡ You must provide exactly one value for each of the following fields: “Country,” “State_or_Province,” and “City”. You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA. Scene Graph: <abstract added scene graph + event agent output + temporal agent output> Prompt: <prompt agent generated geospatial agent prompt>
Temporal Agent Cross Extraction prompt.
You are an agent designed to deduce the time of day, time period, or year depicted in an image based on its visual elements, the scene graph, and a descriptive prompt paying special attention to the added event and geospatial information along with the event’s background. It is important to note that any null, NA, unknown values in the provided are ignored. When analyzing an image, consider the following:
@@
- Best Guess: If conclusive reasoning is not possible, provide the most plausible estimate and explain your reasoning. You will return the output in the required response format. It is absolutely imperative that you return the JSON output.You will extract the information required for the response format. If no plausible guess can be made for a field, output NA. Scene Graph: <abstract added scene graph + event agent output + geospatial agent output> Prompt: <prompt agent generated temporal agent prompt>
Event Agent Cross Extraction prompt.
You are an agent designed to analyze global events using multimodal inputs. You combine the given inputs, paying special attention to the added temporal and spatial information to provide a detailed, concise analysis of the primary event asssociated with this image along with the background behind the primary event. It is important to note that any null, NA, unknown values in the provided temporal and spatial information are ignored. Use logical reasoning and general world knowledge to address gaps in the inputs. Output in the following structure:
@@
3. Best Guess: Make logical assumptions if the inputs are inconclusive, ensuring relevance to the global context. Tailor responses to be clear, neutral, and informative, avoiding bias or unfounded speculation. You will return the output in the required response format. It is absolutely imperative that you return the JSON output.You will extract the information required for the response format. If no plausible guess can be made for a field, output NA. Scene Graph: <abstract added scene graph + geospatial agent output + temporal agent output> Prompt: <prompt agent generated event agent prompt>
Zeroshot COT prompt.
@@
background: the most relative background related to the event
Let’s think step-by-step.You may or may not be provided with the public figures present in the image. If present, ground them to their common nouns in their entities.You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Detective Agent prompt.

### v2

@@ Abstract
Shikhhar Siingh Abhinav Rawat Chitta Baral  Vivek Gupta Arizona State University {ssiingh, arawat11, chitta, vgupt140}@asu.edu
**footnotetext: These authors contributed equally to this work and share first authorship.footnotetext: This author supervised the research and serves as the corresponding author.
@@ 1 Introduction
Motivation: The challenges above underscore why extracting event-specific context from public event images remains difficult. Existing approaches either limit themselves to shallow visual cues or depend on external retrieval, which risks hallucinations, misinformation, and fragmented context. True event understanding requires reasoning about temporal, spatial, and sociopolitical factors—often implicit rather than visually explicit.
@@
Main Contributions: Our approach is driven by the need for a systematic, reasoning-based method to extract three key components from an image: (1) geospatial location (e.g., "Indonesia"), (2) temporal information (e.g., "October 20, 2019"), and (3) event-specific details (e.g., "inauguration of President Joko Widodo’s second term amid public tensions"). To overcome prior limitations, we introduce a structured, multi-agentic framework where specialized agents collaboratively extract and validate contextual elements. Inspired by Ng et al. (2024), this architecture reduces errors and hallucinations through cross-agent validation, improving both precision and reliability. A thorough working example for the framework is explained in Appendix B under Working Example.
@@
The code and dataset are available at https://coral-lab-asu.github.io/getreason/
@@ 2 Dataset Description
The TARA dataset comprises of approximately 16,000 images associated with news articles, along with their corresponding time and location information, automatically extracted from the New York Times (NYT)https://developer.nytimes.com/docs/archive-product/1/overview. For our evaluation, we utilized the train set which comprises of 12,306 samples from this dataset, spanning from January 2010 to May 2021. Each image is accompanied by metadata detailing the time and location pertinent to the associated news article.
@@ 2.1 TARA*
To better suit our task, we augment TARA dataset with event and fine-grained temporal and geospatial information generated from the metadata present in the ground truth using a Vision Language Model (VLM), specifically, Gemini 1.5 Prohttps://deepmind.google/.
@@ 5.1 Experimental Settings
For GPT-4o-minihttps://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, we utilized the batch API. While OpenAI’s models do not directly reveal public figureshttps://openai.com/policies/usage-policies/, they can ground them if their identities are provided externally in the prompt. To leverage this, we used the CelebrityFaces API from AWS Rekognitionhttps://aws.amazon.com/rekognition/ to identify public figures in the dataset images with a confidence score of at least 99%. These identified public figures were then explicitly included in the prompts during the experiments to enable more precise entity recognition.
@@
Baselines.
We used multiple visual reasoning approaches to compare with GETReason. These include: reasoning-based methods such as GoodGuesser (Zhang et al., 2024) and CogBench (Song et al., 2024), which highlight limitations in spatio-temporal inference due to unstructured representations; object detection-based methodologies such as CAPTURE (Dong et al., 2024), VIPHY (Singh et al., 2023), and EDIS (Liu et al., 2023), which focus on descriptive features and offer limited contextual reasoning; and retrieval-augmented generation (RAG) methods like Generate the Select (Fu et al., 2023) and QR-CLIP (Gao and Zhou, 2024), which enhance factual recall but struggle with abstract or causal reasoning. To ensure fairness, all baselines were evaluated under a unified schema using only the VLM’s internal knowledge. Additionally, we introduced two custom-designed baselines tailored to our task, Zeroshot COT and Detective Agent, whose prompts are provided in Appendix A.10 and Appendix A.11.
Evaluation Metric.
@@ 5.6 Effectiveness Analysis
1. Fundamental methodological differences:
a. Hierarchical Multi-Agent Framework:
Our method employs a hierarchical multi-agent framework, reducing cognitive load on the VLM, minimizing hallucinations, and enabling context resetting—unlike baselines that rely on single, monolithic prompts.
b. Structured Prompting:
We use structured, role-specific prompts and controlled information flow, providing greater consistency and interpretability than baselines that depend on ad-hoc prompt engineering.
c. Schema-Driven Reasoning:
By combining structured schema extraction with flexible reasoning, our approach achieves both precision and adaptability, addressing the inefficiencies and hallucinations often seen in less structured methods.
d. Systematic Cue Integration:
Our framework explicitly handles temporal, geospatial, and event-specific cues, which are often overlooked or inconsistently extracted in prior methods (see Section 6).
e. Robust, Large-Scale Evaluation:
GETReason is evaluated on over 17,000 examples, enabling a more robust and generalizable assessment than prior works limited to datasets of around 2,000 samples.
2. Diverse Baseline Limitations:
As detailed in Section 6, baseline methods span object detection, shallow reasoning, and retrieval-based approaches—each introducing specific limitations in contextual depth, prompt precision, or overreliance on external sources.
6 Comparison with Related Work
Reasoning-Based Approaches.
Object Detection-Based Approaches.
RAG-Based Approaches.
@@ 7 Conclusion
The main limitation of this approach is the language barrier, as it relies on English-language datasets. Additionally, the TARA dataset’s extraction method results in many images being unrelated to their corresponding articles. Our analysis is based on a subset of TARA, whereas a more comprehensive evaluation could be conducted on the full 61k dataset or even larger alternatives like SHERLOCK Hessel, Hwang, Park, Zellers, Bhagavatula, Rohrbach, Saenko, and Choi (2022) and BreakingNews dataset Ramisa, Yan, Moreno-Noguer, and Mikolajczyk (2017) along with the comparison with newer studies like PuzzleGPT Ayyubi, Feng, Liu, Lin, Wang, and Chang (2025) to better assess data loss and the study’s limitations. Furthermore, as context extraction depends on model reasoning and Vision-Language Models (VLMs) are rapidly advancing, our findings may not fully reflect the latest developments. Although we used three competitive models, more robust reasoning based alternatives like OpenAI-o1https://openai.com/o1/ , OpenAI-o3-minihttps://openai.com/index/openai-o3-mini/ and DeepSeek-R1 Team (2025a) have since emerged.
@@ Acknowledgements
We thank the anonymous reviewers for their helpful feedback. We also thank Zenia Shaikh for her valuable assistance and contributions to this work. We thank the Complex Data Analysis and Reasoning Lab at Arizona State University for computational support, and Krishna Singh Rajput for his support during the submission process.
Appendix A Individual Agentic Prompts:
A.1 Scene Graph Agent prompt.
You are an expert in analyzing and understanding visual images to generate detailed and structured scene graphs in JSON-formatted outputs. You perform the following two tasks sequentially:
Scene Graph Generation: For every image provided, identify all entities, objects, attributes, and relationships in the scene. The graph must include all common entities, whether singular or plural, such as "car," "person," "crowd," "tree," "building," etc., as well as their properties and the spatial or contextual relationships among them. Your goal is to create a complete and accurate representation of all visible elements in the scene.
Graph Enhancement: Refine the scene graph by identifying and replacing the generic descriptions of entities that correspond to public figures provided in the input list. For instance, if a "person" in the scene is identified as a specific public figure, replace "person" with their provided name. During this process, ensure that:
All entities, whether identified or unidentified, remain in the graph.
All relationships between objects and entities are preserved.
Unidentified or unrecognized entities are left unchanged but still included in the graph.
Your responses must be concise but thorough, breaking down the image into its components and interactions while ensuring that the scene graph remains fully intact, comprehensive, and user-friendly. Do not prune out entities or relationships during any part of the process.
The objective is to provide users with a refined scene graph that includes all visible entities and incorporates specific public figure names where applicable, enhancing both clarity and usability.
You may or may not be provided with the public figures present in the image. If present, ground them to their common nouns in their entities.
You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
A.2 Abstract Agent prompt.
@@ Appendix A Individual Agentic Prompts:
[…]
@@
You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Scene Graph: <generated scene graph>
A.3 Prompt Agent Prompt.
@@
Geospatial specialist - Analyzes locations from images and scene graphs with detailed reasoning. You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Scene Graph:<abstract added scene graph>
A.4 Geospatial Agent Direct Extraction prompt.
You are a custom agent designed to analyze an image and any accompanying contextual information to deduce geospatial details about the scene depicted. The details to determine are:
- Country
- State_or_Province
- City
Using the visual elements of the image and the provided contextual information, you must determine the best possible value for each of these fields. If you cannot deduce a particular field with reasonable confidence, output "NA" for that field.
Rules:
- You must provide exactly one value for each of the following fields: "Country," "State_or_Province," and "City." You will return the output in the required response format. It is absolutely imperative that you return the JSON output.You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.",
Scene Graph: <abstract added scene graph>
Prompt: <prompt agent generated geospatial agent prompt>"
A.5 Temporal Agent Direct Extraction prompt.
You are an agent designed to deduce the time of day, time period, or year depicted in an image based on its visual elements, the scene graph, and a descriptive prompt.
When analyzing an image, consider the following:
@@
- Best Guess: If conclusive reasoning is not possible, provide the most plausible estimate and explain your reasoning. You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Scene Graph: <abstract added scene graph>
Prompt: <prompt agent generated temporal agent prompt>
A.6 Event Agent Direct Extraction prompt.
@@
3. Best Guess: Make logical assumptions if the inputs are inconclusive, ensuring relevance to the global context. Tailor responses to be clear, neutral, and informative, avoiding bias or unfounded speculation.You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Scene Graph: <abstract added scene graph>
Prompt: <prompt agent generated event agent prompt>
A.7 Geospatial Agent Cross Extraction prompt.
You are a custom agent designed to analyze an image and any accompanying contextual information to deduce geospatial details about the scene depicted, paying special attention to the added event and temporal information along with the events background. It is important to note that any null, NA, unknown values in the provided are ignored. The details to determine are:
- Country
- State_or_Province
- City
Using the visual elements of the image and the provided contextual information, you must determine the best possible value for each of these fields. If you cannot deduce a particular field with reasonable confidence, output "NA" for that field.
@@
- You must provide exactly one value for each of the following fields: "Country," "State_or_Province," and "City". You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Scene Graph: <abstract added scene graph + event agent output + temporal agent output>
Prompt: <prompt agent generated geospatial agent prompt>
A.8 Temporal Agent Cross Extraction prompt.
You are an agent designed to deduce the time of day, time period, or year depicted in an image based on its visual elements, the scene graph, and a descriptive prompt paying special attention to the added event and geospatial information along with the events background. It is important to note that any null, NA, unknown values in the provided are ignored. When analyzing an image, consider the following:
@@
- Best Guess: If conclusive reasoning is not possible, provide the most plausible estimate and explain your reasoning. You will return the output in the required response format. It is absolutely imperative that you return the JSON output.You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Scene Graph: <abstract added scene graph + event agent output + geospatial agent output>
Prompt: <prompt agent generated temporal agent prompt>
A.9 Event Agent Cross Extraction prompt.
You are an agent designed to analyze global events using multimodal inputs. You combine the given inputs, paying special attention to the added temporal and spatial information to provide a detailed, concise analysis of the primary event associated with this image along with the background behind the primary event.
It is important to note that any null, NA, or unknown values in the provided temporal and spatial information are ignored. Use logical reasoning and general world knowledge to address gaps in the inputs.
Output in the following structure:
@@
3. Best Guess: Make logical assumptions if the inputs are inconclusive, ensuring relevance to the global context.
Tailor responses to be clear, neutral, and informative, avoiding bias or unfounded speculation. You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
Scene Graph: <abstract added scene graph + geospatial agent output + temporal agent output>
Prompt: <prompt agent generated event agent prompt>
A.10 Zeroshot COT prompt.
@@
background: the most relevant background related to the event
Lets think step-by-step. You may or may not be provided with the public figures present in the image. If present, ground them to their common nouns in their entities. You will return the output in the required response format. It is absolutely imperative that you return the JSON output. You will extract the information required for the response format. If no plausible guess can be made for a field, output NA.
A.11 Detective Agent prompt.